Greta, an elderly investor, has a degree of risk aversion of A= 5 when applied to return on wealth over a one-year horizon. Sshow in excel!
Answer #1

S&P Portfolio Hedge FundPortfolio Risk premiums 0.05 0.12 SDs 0.20 0.40 Sharpe ratios 0.2500 =0.05/0.2 0.3000 =0.12/0.4

*Please rate thumbs up